OpenStudy (anonymous):

Please Help, MEDAL WILL BE REWARDED! Find dy/dx by implicit differentiation and evaluate the derivative at the given point. x^3+y^3=4xy+1, (2,1)

hartnn (hartnn):

so whats derivative of x^3 ? y^3 ?

OpenStudy (anonymous):

its 3y^2 and then 3x^2

hartnn (hartnn):

to get d/dx (y^3) you need to use chain rule know what that is ?

OpenStudy (anonymous):

yes, so I would get 3y^2 again

hartnn (hartnn):

no... d/dx (y^3) = 3y^2 dy/dx

OpenStudy (anonymous):

okay so do I plug in my numbers from there and solve?

hartnn (hartnn):

on the right side you need product rule

OpenStudy (anonymous):

okay hold on...I get 9y^2/x

hartnn (hartnn):

wait how ? on right, d/dx (xy) = x dy/dx + y

OpenStudy (anonymous):

Oh, okay, so I would get -3/x^2

hartnn (hartnn):

where does that come from ? :O i am lost

OpenStudy (anonymous):

Well im using product rule, right? Or am I using it wrong?

OpenStudy (anonymous):

If I use product rule d/dd I get 0

hartnn (hartnn):

ok, let me take you term by term d/dx x^3 = 3x^2 d/dx y^3 = 3y^2 dy/dx d/dx 4xy = 4 (x dy/dx + y) d/dx 1 = 0

OpenStudy (anonymous):

okay so that's d/dx but I need dy/dx, so would I substitute this into my original equation?

hartnn (hartnn):

yes, we need dy/dx when x=2, y=1 so keep dy/dx as it is only and plug in values for x and y

OpenStudy (anonymous):

okay...I would start 2^3+1^3=4(2)(1)+1 right?

hartnn (hartnn):

where did dy/dx go ?

OpenStudy (anonymous):

IDK im confused on what equation im using

hartnn (hartnn):

3x^2 + 3y^2 dy/dx = 4 (x dy/dx + y) x=2 y=1

OpenStudy (anonymous):

okay so my final answer would be 3x^2+3y^2 d(1)/d(2)=4( 2 d(1)/d(2) +1)?? @hartnn

hartnn (hartnn):

no... i said keep dy/dx as it is

hartnn (hartnn):

3 (2^2) +3(1^2) dy/dx = 4 (2* dy/dx+1) find dy/dx from here

OpenStudy (anonymous):

okay, I get 12+4 dy/dx=4(2dy/dx+1) which is 16 dy/dx =4(3dy/dx)

hartnn (hartnn):

no... 12 +3 dy/dx = 8 dy/dx +4 12 -4 = (8-3) dy/dx 8 = 5 dy/dx so. dy/dx = 8/5

OpenStudy (anonymous):

oh I see what your did, I skipped a step in my calculations. Okay, thank you so much! Im going to work this out again and see if I can get it right without making mistakes! Thank you so much!!!

hartnn (hartnn):

good! welcome ^_^